The measure of angle 1 is 4 times the measure of its supplement. What is the measure, in degrees, of angle 1?
Answer by addingup(3677)

Two angles are supplementary when they add up to 180 degrees. Therefore, if the measure of angle 1 is 4 times the measure of its supplement:
Let's call angle 1 4x and its supplement x.
4x + x= 180
5x= 180
x= 180/5
x= 36 is the measure of the supplement.
Since angle 1 is 4 times the measure of its supplement:
36*4= 144 is the measurement of angle 1.